OpenStudy (anonymous):

The length of a rectangle is twice its width. If the area of the rectangle is 162 m2 , find its perimeter.

OpenStudy (anonymous):

you can assume the length to be L and the width to be W so from your words L=2W and the rectangle area is length*width =L*W=2W*W=2W^2=162 ....now you can find W and L